a guitar string is 0.620 m long, and oscillates at 234 Hz. If a player uses his finger to shorten the string to 0.480 m, what is the new frequency?
Answer #1

We know frequency n=1/2l √(T/m)

Where l is the length of the wire

T is the tension And m is the mass of the wire.

[we will assume that √(T/m) will be constant in both case.]

Therefore √(T/m) =n*2l =234*2*0.620 =290.16

New frequency is N1 =1√(T/m)/2l1 = 1*290.16/(2*0.480)= 302.25 Hz